Assuming that the samples come from normal distributions, find the margin of error E given the following.




2. n=150, x-=98, s=9.90, 95% confidence.



3. n=350, x-=125, s=11.18, 99% confidence.



4. n=500, x-=236, s=15.36, 90% confidence.



5. n=785, x-=459, s=21.42, 99% confidence.

1
Expert's answer
2022-05-05T10:35:36-0400

2. The critical value for "\alpha = 0.05,df=n-1=149" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.976013."

Margin of Error: "E=t_c\times\dfrac{s}{\sqrt{n}}=1.976013 \times \dfrac{9.9}{\sqrt{150}}= 1.5973"


3. The critical value for "\alpha = 0.01,df=n-1=349" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.58999."

Margin of Error: "E=t_c\times\dfrac{s}{\sqrt{n}}=2.58999\times \dfrac{11.18}{\sqrt{350}}= 1.5478"


4. The critical value for "\alpha = 0.10,df=n-1=499" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.647913."

Margin of Error: "E=t_c\times\dfrac{s}{\sqrt{n}}=1.647913 \times \dfrac{15.36}{\sqrt{500}}=1.1320"


5. The critical value for "\alpha = 0.01,df=n-1=784" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.582115."

Margin of Error: "E=t_c\times\dfrac{s}{\sqrt{n}}=2.582115 \times \dfrac{21.42}{\sqrt{785}}= 1.9741"